A 0.30-kg block of wood is suspended on a spring.
Answer #1

from the Hook's law

mg = kx

0.3 ( 9.8) = k (2 * 10^-2)

K = 147 N/m

period of spring mass system is

T =2pi sqrt m/k

= 2pi sqrt 0.3/147

=0.2837 s

for 3 comple cycles of vibration

T = 3 (T) = 3 ( 0.2837) = 0.8511 s

----------------------------------------------------------------------

(b)

E = 1/2 * K A^2

= 1/2 * 147 ( 1.0 * 10^-2)^2

=0.00735 J